OpenStudy (anonymous):

Given the function f(x) = 6x2 −13, what is f(-3)?

OpenStudy (lgbasallote):

substitute every x with -3 f(-3) = 6(-3)^2 - 13

OpenStudy (theviper):

i.e., f(-3)=6*9-13=54-13=41